OpenStudy (anonymous):

790(x-12)-25=120 what dose x mean

OpenStudy (angela210793):

790X-790x(-12)=120+25 790X=145-9480 790X=9335 X=9335/790 X=1867/158 Was this what u were looking 4?

OpenStudy (anonymous):

thx umm not realy but u did do a good job

OpenStudy (amistre64):

x means "do not enter"; "danger"; "keep away"... lemming diving board ahead

OpenStudy (angela210793):

U weren't? what are you looking 4 then? O.o

OpenStudy (anonymous):

Please give more info because to find what x means we would need to know whether we are to solve the equation or find what x represents from a word problem and to do that we need the word problem. can you explain the question?